Reports 1-1 of 1 Clear search Modify search
Detector Operation (Operations Report)
magazzu - 0:01 Wednesday 23 January 2019 (44485) Print this report
Operator Report - Afternoon Shift 22 January 2019

The shift was dedicated to the continuation of the recovery of the ITF, carried out by A. Allocca. At around 16:00 UTC we started suffering from some instability of the Input Mode Cleaner, A. Chiummo investigated the issue. Also, at 19:30 UTC, the suspensions started to be unstable due to an earthquake ( see attached image ).

The highest step reached stably was LOCKED_OMC1_B1S2_DC. Activity still in progress.

Images attached to this report
Search
×

Warning

×